OpenStudy (anonymous):

Consider the differential equation y′′+αy′+βy=t+e^(4t). Suppose the form of the particular solution to this differential equation as prescribed by the method of undetermined coefficients is yp(t)=A1t^2+A0t+B0te^(4t). Determine the constants α and β.

OpenStudy (anonymous):

>>

OpenStudy (anonymous):

Derivatives of the particular solution: \[\begin{align*} y_p(t)&=A_1t^2+A_0t+B_0te^{4t}\\\\ {y_p}'(t)&=2A_1t+A_0+B_0(1+4t)e^{4t}\\\\ {y_p}''(t)&=2A_1+B_0(8+16t)e^{4t} \end{align*}\] Subbing into the ODE and matching up coefficients gives the system \[\begin{cases} 2A_1+\alpha A_0=0&\text{(constant)}\\ 2\alpha A_1+\beta A_0=1&(t)\\ \beta A_1=0&(t^2)\\ (8+\alpha)B_0=1&(e^{4t})\\ (16+4\alpha+\beta)B_0=0&(te^{4t}) \end{cases}\] I haven't taken the time to look at this much, but I'm not sure it's guaranteed that you can find numerical values for \(\alpha\) and \(\beta\) unless you're given values for \(A_1,A_0,B_0\)...

OpenStudy (anonymous):

I looked it up online and found that \[\beta = 0\] I would like some help in just how to start solving this problem. I look at it and I don't even know where to start.

OpenStudy (anonymous):

If we assume that \(A_0\neq0\), then the third equation of the system tells us that \(\beta=0\). With this info you can solve for \(\alpha\) in the fifth equation: \[(16+4\alpha)B_0=0~~\implies~~16+4\alpha=0\] (provided that \(B_0\neq0\)).